# A Tutorial on MIMO-OFDM ISAC: From Far-Field to Near-Field

### v1

@@ IV-B3 One-dimensional Parameter Estimation Algorithms
In the previous subsections, we have modeled the one-dimensional (1D) parameter estimation problems in (22), (25), (29), (38) for AoA, delay, and Doppler estimation. It is not difficult to see that all these problems share a similar structure, and thus the problems can be solved using essentially the same algorithms. In this section, we provide examples of 1D AoA estimation algorithms, while they can be easily modified for delay and Doppler estimation.
@@
×$\qquad\qquad\quad\times\sum_{m=0}^{M_{r}-1}e^{j2\pi m\frac{d}{\lambda}(\sin{\theta}-\sin{\theta_{k}})}\bigg{|}^{2}$ (39)
@@
×$\qquad\qquad\quad\times{\frac{{\sin\left({\pi\frac{{d}}{\lambda}M_{r}(\sin\theta-\sin{\theta_{k}})}\right)}}{{\sin\left({\pi\frac{{d}}{\lambda}(\sin\theta-\sin{\theta_{k}})}\right)}}}\bigg{|}^{2}$
$\mathop{\approx}\limits^{(a)}\frac{{{\left|{b_{n,p}}\right|}^{2}}}{M_{r}}\sum_{k=1}^{K}{{{\left|{{\overline{\alpha}}_{k}}\right|}^{2}}}{{{\left|{\frac{{\sin\left({\pi\frac{{d}}{\lambda}M_{r}(\sin\theta-\sin{\theta_{k}})}\right)}}{{\sin\left({\pi\frac{{d}}{\lambda}(\sin\theta-\sin{\theta_{k}})}\right)}}}\right|}^{2}}}$
where $\mathbf{a}_{\mathrm{R}}(\theta)$ is the steering vector in the spatial domain in (18), $\theta\in\left[-90^{\circ},90^{\circ}\right]$ is the observation angle, and $\theta_{k}$ is the $k$-th target angle. The approximation in $(a)$ holds if the angular difference between the targets is larger than the angular resolution of the array. The angular resolution of an array is based on the Rayleigh limit [112], which corresponds to half the width between the null points of a focused beam of the array. Form (IV-B3), the null points for a beam pointed at $\theta_{k}$ can be obtained by $\pi\frac{{d}}{\lambda}M_{r}(\sin{\theta}-\sin{\theta_{k}})=\pm\pi$, so the beamwidth is given by
@@
where $D=M_{r}d$ denotes the array aperture. The maximum resolution is achieved when $\theta_{k}\rightarrow 0^{\circ}$, and then
@@
Thus, the angular resolution of a ULA is inversely proportional to the effective array aperture. This means that as the effective array aperture increases, the beamwidth becomes narrower, thus enhancing the ability of the array to resolve different targets.
Similarly, for delay and Doppler estimation, the power spectrum in the delay and Doppler domains can be derived from signals $\mathbf{X}_{\tau}$ and $\mathbf{X}_{\upsilon}$ in (25) and (29), respectively. By calculating the corresponding null-to-null bandwidth, the minimum resolvable difference in the target delay and Doppler frequency are given by
$\Delta\tau=\frac{{1}}{N\Delta f}=\frac{1}{{B}},\ \Delta\upsilon=\frac{1}{PT_{s}}=\frac{1}{\mathrm{CPI}}.$ (42)
Thus, the delay and Doppler resolution are inversely proportional to system bandwidth $B$ and CPI, respectively. In contrast to the angular resolution, which depends on the target angle, the delay and Doppler resolution are independent of the target parameters.
In practice, calculating the power spectrum in the angular domain (see (IV-B3)) involves an exhaustive search over $\theta\in[-90^{\circ},90^{\circ}]$, resulting in high computational complexity. To address this issue, the periodogram algorithm discretizes the phase $\varphi_{\theta}=\frac{d}{\lambda}\sin{\theta}\in\left[-\frac{d}{\lambda},\frac{d}{\lambda}\right]$ in the angular domain into $N_{\theta}^{\mathrm{IFFT}}$ points and employs the IFFT to accelerate the computation. When $N_{\theta}^{\mathrm{IFFT}}>M_{r}$, zero-padding is utilized to convert $\mathbf{X}_{\theta}\in\mathbb{C}^{M_{r}\times Q_{N,P}}$ in (22) to $\mathbf{X}_{\theta}^{\mathrm{ZP}}\in\mathbb{C}^{N_{\theta}^{\mathrm{IFFT}}\times Q_{N,P}}$. In particular, the periodogram $\mathbf{f}_{\theta}\in\mathbb{C}^{N_{\theta}^{\mathrm{IFFT}}\times 1}$ can be obtained by performing an IFFT on each column of matrix $\mathbf{X}_{\theta}^{\mathrm{ZP}}$ and then applying column-wise summation, which is given by
@@
where ||$\cdot$|| ⋅ | denotes element-wise modulus, and $\mathbf{W}_{N}$ is the $N$-point IDFT matrix, expressed as
@@
with $W=e^{-j\frac{2\pi}{N}}$. Additionally, since the default phase of the IFFT is $[0,\pi]$, a period transformation for $\mathbf{f}_{\theta}$ is required to shift the period to $[-\pi/2,\pi/2]$. Then, the peak index $\hat{n}_{\theta}^{k}$ of the periodogram $\mathbf{f}_{\theta}\in\mathbb{C}^{N_{\theta}^{\mathrm{IFFT}}\times 1}$ indicates the phase $\frac{d}{\lambda}\sin{\hat{\theta}_{k}}=\frac{\hat{n}_{\theta}^{k}}{N_{\theta}^{\mathrm{IFFT}}},\hat{n}_{\theta}^{k}\in\left[-\frac{N_{\theta}^{\mathrm{IFFT}}}{2},\frac{N_{\theta}^{\mathrm{IFFT}}}{2}\right]$, in the angular domain. As such, the AoAs of the sensing targets can be estimated as
@@
TABLE III: Simulation setup for monostatic MIMO-OFDM ISAC.
@@
Radial velocity $(8,12,20)\ \mathrm{m/s}$
Doppler frequency $(1.49,2.24,3.73)\ \mathrm{kHz}$
Consider a monostatic ISAC with the simulation setup given in Table III. The periodogram spectrum for AoA estimation after normalization and conversion to dB is shown in Fig. 6. It is observed that higher spectral peaks occur at groundtruth AoAs, while lower sidelobes are present at other angles. The height of the spectral peaks correspond to the strength of the target echos, and targets with weaker echo signals may be obscured by the sidelobes of stronger targets. When the target angle approaches $0^{\circ}$, the width of the spectral peak becomes narrower, which indicates that it can be more easily distinguished by the periodogram method.
@@
MUSIC: The MUSIC algorithm exploits the orthogonality between the noise and signal subspaces to estimate AoAs/delays/Doppler frequencies. These subspaces can be obtained via an eigenvalue decomposition (EVD) of the covariance matrix of $\mathbf{X}_{\theta}$ in (22). This can be expressed mathematically as
@@
where $\mathbf{R}_{s}=\frac{1}{NP}\mathbf{S}_{\theta}\mathbf{S}^{H}_{\theta}$, $\sigma^{2}$ is the noise power, $\mathbf{\Sigma}_{s}$ and $\mathbf{\Sigma}_{n}$ are diagonal matrices composed of the $K$ largest eigenvalues and the remaining $M_{r}-K$ eigenvalues, respectively, $\mathbf{E}_{s}$ and $\mathbf{E}_{n}$ denote the eigenvectors that span the signal and noise subspaces. As such, the MUSIC spectrum can be written as
@@
For the simulation setup in Table III, the AoA spectrum of the MUSIC algorithm after normalization is shown in Fig. 7. It is observed that both the MUSIC and periodogram algorithms can accurately estimate the AoAs of all three targets. However, the spectrum of MUSIC and periodogram differ primarily in two aspects. First, the spectral peaks of MUSIC are sharper, leading to higher angular resolution. Second, the spectrum of MUSIC is smoother than the periodogram, which can prevent weaker targets from being obscured by the sidelobes of stronger targets, hence reducing estimation errors. It is also important to note that subspace methods such as MUSIC and ESPRIT (discussed next) rely on knowledge of the number of AoAs in order to work properly. This is a drawback compared with the non-parametric periodogram approach.
Figure 7: MUSIC spectrum for AoA estimation in MIMO-OFDM ISAC.
ESPRIT: The ESPRIT algorithm exploits the rotational invariance between two subarrays in an array to estimate AoAs. For the signal model in (22), by taking the first $M_{r}-1$ antennas of the ULA as subarray 1, and the last $M_{r}-1$ antennas as subarray 2, the received signals of the two subarrays can be expressed as
@@
where $\mathbf{A}_{1}=\mathbf{A}_{\theta}(1:M_{r}-1,:)$ and $\mathbf{A}_{2}=\mathbf{A}_{\theta}(2:M_{r},:)$ denote the steering matrices of the two subarrays, respectively, $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ are corresponding noises. The two subarray steering matrices are related by $\mathbf{A}_{2}=\mathbf{A}_{1}\mathbf{\Phi}$, where $\mathbf{\Phi}$ denotes the rotation operator determined by the target angles, and is expressed as
@@
Thus, the target AoAs can be estimated by solving the rotation operator.
For ease of calculation, the data from the two subarrays are concatenated into a new data matrix, i.e.,
@@
Then, the signal subspace $\mathbf{E}_{s}$ and noise subspace $\mathbf{E}_{n}$ can be obtained by performing an EVD on the covariance matrix of the data matrix $\bar{\mathbf{X}}$, given by
$\mathbf{R}_{\bar{\mathbf{X}}}=\frac{1}{NP}\bar{\mathbf{X}}\bar{\mathbf{X}}^{H}=\mathbf{E}_{s}\mathbf{\Lambda}_{s}\mathbf{E}_{s}^{H}+\mathbf{E}_{n}\mathbf{\Lambda}_{n}\mathbf{E}_{n}^{H}.$ (51)
Due to the rotational invariance between the two subarrays, the signal subspace can be decomposed as
$\mathbf{E}_{s}=\begin{bmatrix}\mathbf{E}_{s1}\\ \mathbf{E}_{s2}\end{bmatrix}=\begin{bmatrix}\mathbf{A}_{1}\mathbf{T}\\ \mathbf{A}_{1}\mathbf{\Phi}\mathbf{T}\end{bmatrix},$ (52)
@@
OMP: The OMP algorithm is an efficient sparse recovery method based on iteratively selecting the basis vectors most correlated with the received signal to estimate target parameters, thereby avoiding complex optimization procedures. For AoA estimation, by discretizing observation angle $\theta\in[-90^{\circ},90^{\circ}]$ into $N^{s}$ points, the steering dictionary matrix is defined as
@@
where each spatial steering vector $\mathbf{a}_{\mathrm{R}}(\theta)$ serves as an atom. Then, the matrix $\mathbf{X}_{\theta}$ in (22) is reduced to
$\mathbf{X}_{\theta}=\mathbf{A}_{\theta}\mathbf{S}_{\theta}+\mathbf{Z}_{\theta}=\mathbf{A}_{\mathrm{dic}}\mathbf{\Lambda}\mathbf{S}_{\theta}+\mathbf{Z}_{\theta},$ (57)
where $\mathbf{\Lambda}$ is the sparse representation basis. For any snapshot $q$ of $\mathbf{S}_{\theta}$, the sparse recovery problem can be formulated as
$\min_{\mathbf{S}_{\theta}(:,q)}\|\mathbf{S}_{\theta}(:,q)\|_{0}$ (58)
$\mathrm{s.t.}\ \ \mathbf{X}_{\theta}(:,q)=\mathbf{A}_{\mathrm{dic}}\mathbf{\Lambda}\mathbf{S}_{\theta}(:,q)+\mathbf{Z}_{\theta}(:,q),$
where $\|\cdot\|_{0}$ denotes the $\ell_{0}$ norm. Since $\ell_{0}$-norm minimization is NP-hard, greedy optimization algorithms are typically employed to approximate the original signal via a sparse representation. The OMP algorithm iteratively selects the most correlated atoms from the discretized dictionary, updates the residual via least squares minimization, and ultimately estimates the angles of $K$ targets.
Specifically, OMP first initializes the residual as the signal $\mathbf{X}_{\theta}$ in (22) and the support set as the empty set $\emptyset$, i.e.,
$\mathbf{R}_{0}=\mathbf{X}_{\theta},\mathbb{S}_{\hat{\theta}}=\emptyset.$ (59)
The algorithm then performs $K$ iterations, each selecting the atom most correlated with the current residual, which can be expressed by
@@
and then the support set is updated as $\mathbb{S}_{\hat{\theta}}=\mathbb{S}_{\hat{\theta}}\cup\{i_{k}\}$. Here, $\|\cdot\|_{2}$ is the Frobenius norm and $\mathbf{R}_{k}$ the residual after the $k$-th iteration, where the sparse recovery signal is estimated via least squares as
$\hat{\mathbf{S}}_{\theta,k}$ $=\mathrm{arg}\max_{\mathbf{S}}\|\mathbf{X}_{\theta}-\mathbf{A}_{\mathrm{dic}}(:,\mathbb{S}_{\hat{\theta}})\mathbf{S}\|_{2}$ (61)
@@
Finally, the AoAs are estimated as $\hat{\mathbf{\theta}}=\theta_{\mathrm{grid}}(\mathbb{S}_{\hat{\theta}})$, where $\theta_{\mathrm{grid}}$ denotes the discretized angular grid.
From (41), the angular resolution of the 16-antenna ULA is $\Delta\theta=7.1^{\circ}$. As a further illustration, we consider another scenario with more closely spaced targets located at $(-5,0,10)^{\circ}$, while the other parameters in Table III remain unchanged. We show a comparison in Fig. 8 between the periodogram, MUSIC, and OMP. It is observed that when the angle difference between targets is smaller than the resolution limit of the array, their spectral peaks blend together in the periodogram, making them indistinguishable. However, the OMP and MUSIC algorithms are still able to resolve all three targets. Unlike MUSIC, although the OMP algorithm can also achieve super-resolution, its estimation accuracy significantly deteriorates due to mutual interference between the closely spaced targets. In general, the OMP algorithm exhibits degraded estimation performance in scenarios where targets are densely distributed. For the OMP algorithm, the parameters of each target are estimated separately, preventing weak targets from being masked by strong ones. The sequential one-domain estimation approach in Fig. 4 is fundamentally based on the path-separation idea.
@@
In Section IV-B, we introduced the two decoupled parameter estimation methods and the corresponding 1D estimation algorithms. In Figs. 6, 7 and 8, results for the periodogram, MUSIC, and OMP are presented for AoA estimation in parallel one-domain estimation. In the following, we give an example of using MUSIC for sequential one-domain estimation. For the simulation setup in Table III, we consider $K=5$ targets with AoA, distance, and radial velocity parameters being ($-20^{\circ}$,$20\ \mathrm{m}$,$8\ \mathrm{m/s}$), ($10^{\circ}$,$45\ \mathrm{m}$,$14\ \mathrm{m/s}$), ($10^{\circ}$,$80\ \mathrm{m}$,$20\ \mathrm{m/s}$), $\left(45^{\circ},60\ \mathrm{m},6\ \mathrm{m/s}\right)$, $\left(45^{\circ},45\ \mathrm{m},12\ \mathrm{m/s}\right)$. The corresponding delays and Doppler frequency shifts of the five sensing targets are $\left(0.13\ \mu s,1.49\ \mathrm{kHz}\right)$, $\left(0.3\ \mu s,2.61\ \mathrm{kHz}\right)$, $\left(0.53\ \mu s,3.73\ \mathrm{kHz}\right)$, $\left(0.4\ \mu s,1.12\ \mathrm{kHz}\right)$, $\left(0.3\ \mu s,2.24\ \mathrm{kHz}\right)$, respectively. Under the above setting, the estimation results of sequential one-domain estimation based on MUSIC in the three domains for the target with parameters $\left(10^{\circ},0.3\ \mu s,2.61\ \mathrm{kHz}\right)$ are shown in Fig. 9, where ZF-based beamforming is exploited to separate the target signals with different AoAs and delays. It is observed that for step 1, only three AoAs can be estimated for the five sensing targets. After spatial domain beamforming towards $10^{\circ}$, two delays are estimated in step 2. Finally, after delay domain beamforming for the target with delay $0.3\ \mu s$, the Doppler frequency $2.61\ \mathrm{kHz}$ can be estimated in step 3.

### v2

@@ IV-B3 One-dimensional Parameter Estimation Algorithms
In the previous subsections, we have modeled the one-dimensional (1D) parameter estimation problems in (22), (25), (29), (38) for AoA, delay, and Doppler estimation. It is not difficult to see that all these problems share a similar structure, and thus the problems can be solved using essentially the same algorithms. In this section, we provide examples of 1D AoA estimation algorithms that can be easily modified for delay and Doppler estimation.
@@
$\qquad\qquad\quad\times\sum_{m=0}^{M_{r}-1}e^{j2\pi m\frac{d}{\lambda}(\sin{\theta}-\sin{\theta_{k}})}\bigg|^{2}$ (39)
@@
$\qquad\qquad\quad\times{\frac{{\sin\left({\pi\frac{{d}}{\lambda}M_{r}(\sin\theta-\sin{\theta_{k}})}\right)}}{{\sin\left({\pi\frac{{d}}{\lambda}(\sin\theta-\sin{\theta_{k}})}\right)}}}\bigg|^{2}$
$\mathop{\approx}\limits^{(a)}\frac{{{\left|{b_{n,p}}\right|}^{2}}}{M_{r}}\sum_{k=1}^{K}{{{\left|{{\overline{\alpha}}_{k}}\right|}^{2}}}{{{\left|{\frac{{\sin\left({\pi\frac{{d}}{\lambda}M_{r}(\sin\theta-\sin{\theta_{k}})}\right)}}{{\sin\left({\pi\frac{{d}}{\lambda}(\sin\theta-\sin{\theta_{k}})}\right)}}}\right|}^{2}}},$
where $\mathbf{a}_{\mathrm{R}}(\theta)$ is the steering vector in the spatial domain in (18), $\theta\in\left[-90^{\circ},90^{\circ}\right]$ is the observation angle, and $\theta_{k}$ is the $k$-th target angle. The approximation $(a)$ holds when the angular difference between the targets is larger than the angular resolution of the array. The angular resolution of an array is based on the Rayleigh limit [123], which corresponds to half the width between the null points of a focused beam of the array. From (IV-B3), the null points for a beam pointed at $\theta_{k}$ can be obtained by $\pi\frac{{d}}{\lambda}M_{r}(\sin{\theta}-\sin{\theta_{k}})=\pm\pi$, so the beamwidth is given by
@@
where $D=M_{r}d$ denotes the array aperture. The maximum resolution is achieved when $\theta_{k}\rightarrow 0^{\circ}$, which implies
@@
Thus, the angular resolution of a ULA is inversely proportional to the effective array aperture; as the effective array aperture increases, the beamwidth becomes narrower, thus enhancing the ability of the array to resolve different targets.
Similarly, for delay and Doppler estimation, the power spectrum in the delay and Doppler domains can be derived from signals $\mathbf{X}_{\tau}$ and $\mathbf{X}_{\upsilon}$ in (25) and (29), respectively. By calculating the corresponding null-to-null bandwidth, the minimum resolvable differences in the target delay and Doppler frequency are given by
$\Delta\tau=\frac{{1}}{N\Delta f}=\frac{1}{{B}},\ \Delta\upsilon=\frac{1}{PT_{s}}=\frac{1}{\mathrm{CPI}},$ (42)
respectively. Thus, the delay and Doppler resolution are inversely proportional to system bandwidth $B$ and CPI, respectively. In contrast to angular resolution, which depends on the target angle, the delay and Doppler resolution are independent of the target parameters.
In practice, calculating the power spectrum in the angular domain (see (IV-B3)) involves an exhaustive search over $\theta\in[-90^{\circ},90^{\circ}]$, resulting in high computational complexity. To address this issue, the periodogram algorithm discretizes the phase $\varphi_{\theta}=\frac{d}{\lambda}\sin{\theta}\in\left[-\frac{d}{\lambda},\frac{d}{\lambda}\right]$ in the angular domain into $N_{\theta}^{\mathrm{IFFT}}$ points and employs the IFFT to accelerate the computation. When $N_{\theta}^{\mathrm{IFFT}}>M_{r}$, zero-padding is utilized to convert $\mathbf{X}_{\theta}\in\mathbb{C}^{M_{r}\times Q_{N,P}}$ in (22) to $\mathbf{X}_{\theta}^{\mathrm{ZP}}\in\mathbb{C}^{N_{\theta}^{\mathrm{IFFT}}\times Q_{N,P}}$. In particular, the periodogram $\mathbf{f}_{\theta}\in\mathbb{C}^{N_{\theta}^{\mathrm{IFFT}}\times 1}$ can be obtained by performing an IFFT on each column of matrix $\mathbf{X}_{\theta}^{\mathrm{ZP}}$ and then applying column-wise summation, given by
@@
where $|\cdot|$ denotes element-wise modulus, and $\mathbf{W}_{N}$ is the $N$-point IDFT matrix, expressed as
@@
with $W=e^{-j\frac{2\pi}{N}}$. Additionally, since the default phase of the IFFT is $0,\pi]$, the period of $\mathbf{f}_{\theta}$ must be shifted to $\pi/2,\pi/2]$. Then, the peak index $\hat{n}_{\theta}^{k}$ of the periodogram $\mathbf{f}_{\theta}\in\mathbb{C}^{N_{\theta}^{\mathrm{IFFT}}\times 1}$ indicates the phase $\frac{d}{\lambda}\sin{\hat{\theta}_{k}}=\frac{\hat{n}_{\theta}^{k}}{N_{\theta}^{\mathrm{IFFT}}},\hat{n}_{\theta}^{k}\in\left[-\frac{N_{\theta}^{\mathrm{IFFT}}}{2},\frac{N_{\theta}^{\mathrm{IFFT}}}{2}\right]$, and the AoAs of the sensing targets can be estimated as
@@
TABLE V: Simulation setup for monostatic MIMO-OFDM ISAC.
@@
Radial velocity (8,12,20)m/s(8,12,20)\ $\mathrm${m/s}
Doppler frequency (1.49,2.24,3.73)kHz(1.49,2.24,3.73)\ $\mathrm${kHz}
Consider a monostatic ISAC system with parameters given in Table V. The periodogram spectrum for AoA estimation after normalization and conversion to dB is shown in Fig. 10. It is observed that higher spectral peaks occur at groundtruth AoAs, while lower sidelobes are present at other angles. The height of the spectral peaks corresponds to the strength of the target echoes, and targets with weaker echo signals may be obscured by the sidelobes of stronger targets. When the target angle approaches $0^{\circ}$, the width of the spectral peak becomes narrower, which enables the target to be more easily distinguished.
@@
Figure 11: MUSIC spectrum for AoA estimation in MIMO-OFDM ISAC.
MUSIC: The MUSIC algorithm exploits the orthogonality between the noise and signal subspaces to estimate AoAs/delays/Doppler frequencies. These subspaces can be obtained via an eigenvalue decomposition (EVD) of the covariance matrix of $\mathbf{X}_{\theta}$ in (22), which is expressed as
@@
where $\mathbf{R}_{s}=\frac{1}{NP}\mathbf{S}_{\theta}\mathbf{S}^{H}_{\theta}$, $\sigma^{2}$ is the noise power, $\mathbf{\Sigma}_{s}$ and $\mathbf{\Sigma}_{n}$ are diagonal matrices composed of the $K$ largest eigenvalues and the remaining $M_{r}-K$ eigenvalues, respectively, $\mathbf{E}_{s}$ and $\mathbf{E}_{n}$ denote the eigenvectors that span the signal and noise subspaces. The MUSIC spectrum can be written as
@@
For the simulation setup in Table V, the AoA spectrum of the MUSIC algorithm after normalization is shown in Fig. 11. Both the MUSIC and periodogram algorithms can accurately estimate the AoAs of all three targets, although the MUSIC and periodogram spectra differ in two aspects. First, the spectral peaks of MUSIC are sharper, implying higher angular resolution. Second, the MUSIC spectrum has essentially no sidelobes, allowing for better detection of weak targets. However, subspace methods such as MUSIC and ESPRIT (discussed next) rely on knowledge of the number of AoAs in order to work properly, which is a drawback compared with the non-parametric periodogram approach.
ESPRIT: The ESPRIT algorithm exploits the rotational invariance between two subarrays to estimate AoAs. For the signal model in (22), we can take the first $M_{r}-1$ antennas of the ULA as subarray 1, and the last $M_{r}-1$ antennas as subarray 2, and the signals received by the two subarrays can be expressed as
@@
where $\mathbf{A}_{1}=\mathbf{A}_{\theta}(1:M_{r}-1,:)$ and $\mathbf{A}_{2}=\mathbf{A}_{\theta}(2:M_{r},:)$ denote the steering matrices of the two subarrays, respectively, $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ are the corresponding subarray noise matrices. The two subarray steering matrices are related by $\mathbf{A}_{2}=\mathbf{A}_{1}\mathbf{\Phi}$, where $\mathbf{\Phi}$ denotes a rotation operator determined by the target angles, expressed as
@@
Thus, the target AoAs can be determined by finding the rotation operator.
The data from the two subarrays are concatenated into a new data matrix, i.e.,
@@
The signal subspace $\mathbf{E}_{s}$ and noise subspace $\mathbf{E}_{n}$ can be obtained by performing an EVD on the covariance matrix of the data matrix $\mathbf{X}_{\theta}$}, given by
$\mathbf{R}_{{\mathbf{X}_{\theta}}}=\frac{1}{NP}{\mathbf{X}_{\theta}}{\mathbf{X}_{\theta}}^{H}=\mathbf{E}_{s}\mathbf{\Lambda}_{s}\mathbf{E}_{s}^{H}+\mathbf{E}_{n}\mathbf{\Lambda}_{n}\mathbf{E}_{n}^{H}.$ (51)
Due to the rotational invariance between the two subarrays, the signal subspace $\bar{\mathbf{E}}_{s}$ of the data matrix $\bar{\mathbf{X}}$ can be written as
$\bar{\mathbf{E}}_{s}=\begin{bmatrix}\mathbf{E}_{s1}\\ \mathbf{E}_{s2}\end{bmatrix}=\begin{bmatrix}\mathbf{A}_{1}\mathbf{T}\\ \mathbf{A}_{1}\mathbf{\Phi}\mathbf{T}\end{bmatrix}$, (52)
@@
OMP: The OMP algorithm is an efficient sparse recovery method based on iteratively selecting the basis vectors most correlated with the received signal to estimate target parameters, thereby avoiding complex optimization procedures. For AoA estimation, the observation angle $\theta\in[-90^{\circ},90^{\circ}]$ is discretized into $N^{s}$ points, and the steering dictionary matrix is defined as
@@
where each spatial steering vector $\mathbf{a}_{\mathrm{R}}(\theta)$ serves as an atom. Then, the expression for $\mathbf{X}_{\theta}$ in (22) is modified as
$\mathbf{X}_{\theta}=\mathbf{A}_{\theta}\mathbf{S}_{\theta}+\mathbf{Z}_{\theta}=\mathbf{A}_{\mathrm{dic}}\mathbf{S}_{\theta}^{\prime}+\mathbf{Z}_{\theta},$ (57)
where $\mathbf{S}_{\theta}^{\prime}\in\mathbb{C}^{N^{s}\times Q_{N,P}}$ is the sparse signal to be reconstructed with $K$ nonzero rows. Then, the sparse recovery problem can be formulated as
$\min_{\mathbf{S}_{\theta}^{\prime}}\|\mathbf{S}_{\theta}^{\prime}\|_{2,0}\ \ \mathrm{s.t.}\ \|\mathbf{X}_{\theta}-\mathbf{A}_{\mathrm{dic}}\mathbf{S}_{\theta}^{\prime}\|_{2}<\epsilon\,$ (58)
where $\|\cdot\|_{2,0}$ is the $\ell_{2,0}$ mixed norm that denotes the number of nonzero rows, $\|\cdot\|_{2}$ is the Frobenius norm, $\epsilon$ is a small constant that represents the error tolerance governed by the noise level. Since $\ell_{0}$-norm minimization is NP-hard, greedy optimization algorithms are typically employed to approximate the original signal via a sparse representation. The OMP algorithm iteratively selects the most correlated atoms from the discretized dictionary, updates the residual via least squares minimization, and ultimately estimates the angles of $K$ targets.
In particular, OMP first initializes the residual as the signal $\mathbf{X}_{\theta}$ in (22) and the support set as the empty set $\emptyset$, i.e.,
$\mathbf{R}_{0}=\mathbf{X}_{\theta},\mathbb{S}_{\hat{\theta}}$$\emptyset$. (59)
The algorithm then performs $K$ iterations, each selecting the atom most correlated with the current residual, expressed as
@@
[…]
$\hat{\mathbf{S}}_{\theta,k}$ $=\mathrm{arg}\min_{\mathbf{S}}\|\mathbf{X}_{\theta}-\mathbf{A}_{\mathrm{dic}}(:,\mathbb{S}_{\hat{\theta}})\mathbf{S}\|_{2}$ (61)
@@
[…]
@@
We consider another scenario with more closely spaced targets located at $(-5,0,10)^{\circ}$, while the other parameters in Table V remain unchanged. We show a comparison in Fig. 12 between the periodogram, MUSIC, and OMP. It is observed that when the angle difference between targets is smaller than the resolution limit of the array ($7.1^{\circ}$ for a 16-element ULA), their spectral peaks blend together in the periodogram, making them indistinguishable. However, the OMP and MUSIC algorithms are still able to resolve all three targets. Although the OMP algorithm can also achieve super-resolution, unlike MUSIC its estimation accuracy significantly deteriorates due to high correlation between the steering vectors of the closely spaced targets. In general, the OMP algorithm exhibits degraded estimation performance in scenarios where targets are densely distributed. For the OMP algorithm, the parameters of each target are estimated separately, preventing weak targets from being masked by strong ones.
Figure 13: Parallel one-domain parameter estimation based on the periodogram in MIMO-OFDM ISAC.
In Section IV-B, we introduced two decoupled parameter estimation methods and their corresponding 1D estimation algorithms. In the following, we provide an example comparing parallel and sequential one-domain estimation. For the simulation setup in Table V, we consider $K=5$ targets with the AoA, distance, and radial velocity parameters ($20^{\circ}$,20m20\ $\mathrm${m},8m/s8\ $\mathrm${m/s}), ($10^{\circ}$,45m45\ $\mathrm${m},14m/s14\ $\mathrm${m/s}), ($10^{\circ}$,80m80\ $\mathrm${m},20m/s20\ $\mathrm${m/s}), $\left(45^{\circ},60\ \mathrm{m},6\ \mathrm{m/s}\right)$, $\left(45^{\circ},45\ \mathrm{m},12\ \mathrm{m/s}\right)$. The corresponding delays and Doppler frequency shifts of the five sensing targets are $\left(0.13\ \mu s,1.49\ \mathrm{kHz}\right)$, $\left(0.3\ \mu s,2.61\ \mathrm{kHz}\right)$, $\left(0.53\ \mu s,3.73\ \mathrm{kHz}\right)$, $\left(0.4\ \mu s,1.12\ \mathrm{kHz}\right)$, $\left(0.3\ \mu s,2.24\ \mathrm{kHz}\right)$, respectively. The estimation and grouping results for the angles and delays of the five targets for parallel one-domain estimation are illustrated in Fig. 13, and the results for sequential one-domain estimation in the three domains of the target with parameters $\left(10^{\circ},0.3\ \mu s,2.61\ \mathrm{kHz}\right)$ are shown in Fig. 14, where ZF-based beamforming is exploited to separate the target signals with different AoAs and delays. Using parallel estimation, the 1D periodogram can only resolve three and four targets in the angular and delay domains, respectively. However, after parameter grouping and power detection based on (35), the sensing performance closely follows that of the 2D periodogram. The sequential estimation involves three steps in sensing a target. For step 1, only three AoAs can be estimated for the five sensing targets. After spatial domain beamforming towards $10^{\circ}$, two delays are estimated in step 2. Finally, after delay domain beamforming for the target with delay $0.3\ \mu s$, the Doppler frequency 2.61kHz2.61\ $\mathrm${kHz} can be estimated in step 3.